Warriorz13:

The half-life of 232Th is 1.4x 10^10 years. If there are 25.0 g or the sample left after 2.8 x 10^10 years, how many grams were in the original sample?

Warriorz13:

@JustSaiyan

Warriorz13:

@ThisGirlPretty Just look lol

ThisGirlPretty:

Jesus o.o What kind of teacher would torture such a soul

Warriorz13:

The nicest teacher I got

ThisGirlPretty:

XD I bet, I wish voca was me teacher :I Id understand things better and make better grades :O

Vocaloid:



Warriorz13:

Aaaand... confused...

Vocaloid:

A = final amount = 25.0g A_0 = initial amount (what you are solving for) h = half life = 1.4x 10^10 t = time = 2.8 x 10^10 years plug in these values and solve for A0

JustSaiyan:

2.8 ⋅ 10^10 ÷ 1.4 ⋅ 10^10 = 2 25 ÷ (1/2) = 50 ÷ (1/2) = 100g There were 100 grams in the original samle.

JustSaiyan:

Sample*

Vocaloid:

@Warriorz13 do you understand how that calculation was done?